1. Consider the following probability distribution for stocks A and B: (Hint: Use five decimal places for the numbers in your step by step solutions please. no excel solutions
Answer #1

a. Return of stock A = 8.8%
Return of stock B= 1.2%

b. Variance of stock A = 5.76%
Variance of stock B = 5.76%


C. Standard deviation of stock A = 2.4%
STANDARD DEVIATION OF STOCK B = 2.4%

D. Coefficient of variation (CV)= standard deviation/Return.

CV of A = 2.4/8.8 = 0.27
CV of B = 2.4/1.2 = 2

Coefficient of variation higher the more Risky.
Stock B is more risky.


E. Calculation of expected return of the portfolio:

The weight in the stock A= 0.75
The weight in the stock B= 0.25
Expected return of stock A = 8.8%
Expected return of stock B = 1.2%

The expected return of the portfolio: weighted average return.
0.75 x8.8+0.25 x1.2
=6.9%

F. Calculation of expected return of the portfolio:

The weight in the stock A= Wa
The weight in the stock B= 1-Wa
Expected return of stock A = 8.8%
Expected return of stock B = 1.2%

The expected return of the portfolio: weighted average return.
8= Wa x8.8+(1-Wa)x1.2
Wa= 0.8947

The weight in the stock A= 89.47%
The weight in the stock B= 10.53%

& 0.30 0.60 12 %. 0-10 A 3 PA PB P(A-A)? P (B-8)? 124. – 24. 84. zy. Ox@ Ox® 0 (0-8-8-4) 06-1-2)2 44. 64. Total EPA-A – 8.8%.